### v1

@@ 4 Study Finding
Training bugs: All four approaches perform poorly in localizing training bugs, as shown in Table 13. DNNLOC and BLUiR perform comparatively better than the other techniques. Table 15 [40] shows a training bug in the fast.ai project, where the combined usage of Gradient Accumulation and the MixedPrecision Callback (Fig. 11) (as shown in Appendix A) leads to improperly scaled and artificially high training loss value. DNNLOC retrieved the ground truth file at the Top@4 position for this bug, whereas BLUiR retrieved the ground truth file at the Top@7 position. Although both DNNLOC and BLUiR demonstrate comparable performance in locating training bugs, they perform poorly overall (e.g., MAP 40.3% - 41.1%). Moreover, BugLocator performs the worst (e.g., Top@23 position for the example bug in Table 15) in localizing the training bugs (Fig. 12) (as shown in Appendix A), which suggests that similarity analysis between bug reports and source code might not be sufficient for identifying these bugs. Training bugs, which occur during the model’s training phase and might involve aspects such as an improperly defined loss function, are more conceptual. These bugs might not be easily located using these existing methods and might require techniques that can offer a deeper insight into the model’s architecture and training process.
@@
Tensor bugs: Tensors are central to deep learning and often involve intricate dimensional and mathematical issues. From Fig. 5, we notice that DNNLOC and BLUiR are more effective in localizing tensor bugs than other techniques (e.g., BugLocator, BLIA). In the example bug from Table 16 [41], the bug report described the issue with the ‘nd.slice’ function in MXNet, which should return an empty tensor when the ‘begin’ and ‘end’ parameters are equal. Instead, the function returns a tensor with the same shape as the data. For this example tensor bug, both DNNLOC and BLUiR retrieved the correct buggy file at the top position. Interestingly, by parsing the AST of the code, BLUiR identified the relevant code snippet in the test_slice() function, which shares significant keyword overlap with the bug report. BLUiR determines the relevance of low-level code elements (e.g., class names, method names) against a bug report, which helps to reduce noise in the code segments where tensors could be handled or manipulated.
@@
API bugs: From Fig. 5, we observe that DNNLOC and BugLocator perform well in localizing API bugs, whereas BLUiR is the least effective. This could be due to BLUiR’s heavy reliance on structural information from the source code, which might not capture the specifics of API bugs (e.g., incorrect API calls). Another factor could be the rapid evolution of deep learning APIs, which affects versioning and compatibility [66]. Due to frequent structural changes in the code, the technique might be less effective in assessing the relevance between code and bug reports.
@@
GPU bugs: Our investigation reveals that GPU bugs are the most difficult to localize for all four existing approaches. One possible explanation could be the complex nature of GPU bugs, as they can be triggered by a variety of factors, such as the compatibility between hardware (e.g., GPU device) and software (e.g., PyTorch). It might not even be located in the source code (a.k.a extrinsic GPU bug) [30]. From Table 20, we notice that only 17.65% of the GPU bugs can be found in the source code (a.k.a intrinsic GPU bug). Some examples of intrinsic GPU bugs that we found – the wrong reference to a GPU device, failed parallelism, incorrect state sharing between subprocesses, and faulty data transfer to a GPU device.
Table 18 [43] shows a GPU bug triggered by the codebase (a.k.a intrinsic). The bug is connected to the use of multiple GPUs during training. According to the report, the machine contains multiple GPU devices, but only one GPU is used during computation. All four bug localization techniques performed poorly in locating the actual buggy code for this GPU bug. BLUiR retrieved the buggy file at the 50th position, followed by BugLocator at the 65th, BLIA at the 47th, and DNNLOC at the 43rd position in their respective rankings.
@@
We observed a minimal keyword overlap and structural similarity between the bug report (Table 18) and the actual buggy code (Fig. 16) (as shown in Appendix A), presenting a significant challenge in locating such bugs using BugLocator. Additionally, BLUiR’s analysis of smaller code segment similarity also proves ineffective for GPU bugs, as it solely concentrates on source code and overlooks the hardware-software interactions and runtime specifics crucial for comprehending such bugs. Similarly, BLIA’s integration of stack trace information, commit history, or version control history fails to encapsulate the unique aspects of GPU bugs as well. Moreover, despite DNNLOC’s capability to capture non-linear complex relationships through deep neural networks, it did not help to locate GPU bugs. One reason might be that these bugs often involve complex hardware-software dynamics that are not typically addressed by standard source code analysis or within the codebase’s non-linear mappings.
[…]
Table 19: Experimental result of existing bug localization techniques (BugLocator, BLUiR, BLIA, DNNLOC) for NDL Bugs in deep learning systems
Method MAP MRR BugLocator 0.362 0.417 BLUiR 0.292 0.334 BLIA 0.437 0.381 DNNLOC 0.588 0.439
@@
[…]
Most of the traditional bug localization techniques rely on the similarity between bug reports and source code [11, 6, 12, 17, 14, 13, 23, 48]. However, if a bug is of extrinsic nature (e.g., originates from the operating system), simply relying on source code may not be effective for its localization.
@@
[…]
Prevalence ratio of extrinsic & intrinsic bugs: We found 40.00% extrinsic bugs within a total of 385 bug reports from deep learning systems (Denchmark dataset). The notion of extrinsic bugs is relatively new, especially in the case of bugs from deep learning systems. For a better comparison, we also manually inspected 385 bugs from non-deep learning systems (BugGL dataset) and determined the prevalence ratio of extrinsic and intrinsic bugs. We found only 10.65% extrinsic bugs in non-deep learning systems. Thus, deep learning systems contain almost four times more extrinsic bugs (Fig. 6) than non-deep learning systems.
[…]
@@
Figure 7: Performance of existing approaches (BugLocator, BLUiR, BLIA, DNNLOC) for localizing extrinsic and intrinsic bugs in both systems (DLSW & NDLSW)
[…]
From Table 21, we notice that DNNLOC performs slightly better than other techniques in localizing extrinsic bugs for both systems. However, it shows a clear performance gap in localizing extrinsic and intrinsic bugs. The technique is less effective with extrinsic bugs, particularly in DLSW. It suggests a shortcoming in handling external complexities despite being able to capture non-linear relationships between bug reports and source code. We notice that BugLocator’s localization performance for extrinsic bugs in DLSW is lower than NDLSW. BugLocator might not be able to locate the extrinsic bugs in deep learning systems due to its naive approach, i.e., considering code as regular texts. We also found that BLUiR shows less performance gap between extrinsic and intrinsic bugs. It extracts different structured items, (e.g., methods, classes) from the source code and bug reports. Thus, even if they reside outside the current codebase and are invoked from an external library, they could be matched with relevant keywords from a bug report. Interestingly, BLIA performs slightly better for extrinsic bugs in DLSW than NDLSW, which could be possible due to its diverse use of meta components (e.g., stack traces, version control history). Stack traces from deep learning systems (DLSW) often contain more intricate information and dependencies (e.g., complex neural network data flows, dependencies on specialized libraries, GPU-related synchronization issues) related to the bugs, unlike stack traces from non-deep learning systems, which might not provide the same level of detailed information [68, 69].
Overall, these results suggest that extrinsic bugs are hard to localize, whether related to deep learning or not. However, deep learning bugs with an extrinsic nature are the more difficult to localize. On the other hand, from Fig. 7, we note that the performance of all four approaches for intrinsic bugs in NDLSW is higher compared to the intrinsic bugs in DLSW, which supports the fact that the bugs related to deep learning algorithms (a.k.a DL bugs) from deep learning systems are more challenging to localize.
@@
Method DLSW+EXT DLSW+INT NDLSW+EXT NDLSW+INT MAP BugLocator 0.287 0.359 0.375 0.405 BLUiR 0.277 0.369 0.301 0.359 BLIA 0.321 0.449 0.382 0.487 DNNLOC 0.334 0.486 0.390 0.505 MRR BugLocator 0.371 0.489 0.409 0.476 BLUiR 0.296 0.445 0.309 0.364 BLIA 0.385 0.518 0.404 0.505 DNNLOC 0.398 0.544 0.420 0.521 DLSW+EXT = Extrinsic Bug from Deep Learning Systems
@@ 4.4 Key findings
Based on our findings from three research questions, we discuss the key factors that challenge the bug localization in deep learning systems as follows:
  • Extrinsic nature of bugs: A substantial number of bugs in deep learning systems are extrinsic, arising from external factors (e.g., hardware compatibility, OS issues). These bugs might not be found through source code analysis alone, making them inherently challenging to localize using traditional methods relying on source code. We also found a strong statistical correlation between extrinsic bugs and bugs from deep learning systems.
  • Textual representation limitations in dependencies: Our baseline techniques heavily rely on textual similarity between bug reports and source code, which is less effective for deep learning bugs. They often involve intricate dependencies (e.g., gradient vanishing in RNN) and require conceptual understanding (e.g., improperly defined loss function). These complex issues may not be represented using texts, which makes the existing baseline techniques less effective.
  • Dynamic complexity in training and tensor operations: Training bugs and tensor bugs are challenging to localize due to their dynamic and complex nature, often involving issues in the training process (e.g., misconfigured batch sizes) and data manipulation (e.g., incorrect tensor reshaping) respectively, which are not adequately captured by existing methods for bug localization.
  • Multifaceted extrinsic influences on GPU operations: GPU bugs are particularly challenging to localize because they can be caused by multiple factors, such as hardware-software interaction and compatibility issues, and may not even be present in the source code (extrinsic bugs).
  • Inadequate information for API usage: Traditional techniques (e.g., BLUiR) struggle with locating API bugs in deep learning systems. Deep learning APIs differ from traditional APIs [66] since they handle complex data structures, specialized hardware dependencies, dynamic computation graphs, high-level abstractions for operation, and rapid evolution, which impacts versioning and compatibility. Traditional bug localization techniques might not be able to capture such intricate information adequately from deep learning APIs, which might be hurting the localization process.
  • Ineffectiveness of stack traces for tensor bugs: For tensor and input bugs, stack trace information is found to be less effective since these bugs are often related to data manipulation and computations rather than the execution flow captured by stack traces.
  • Lack of details on model architecture and training process: Although bug reports in deep learning systems contain more code snippets, this does not significantly improve the performance of bug localization. Deep learning bugs require a deeper understanding of the model architecture and training process, which may not be detailed in the bug reports and the code snippets.
  • Limited scope of non-linearity: While leveraging non-linear relationships between bug reports and source code using deep neural networks has improved the overall performance of bug localization, the significant performance gap between deep learning systems and non-deep learning systems indicates the inherent localization challenges in deep learning systems. It might help localize specific deep-learning bugs (e.g., model bugs) but might not others (e.g., training and GPU bugs). It suggests that the deep learning approach, despite its ability to capture non-linear relationships, fails to address the complexities of the DL training process or the GPU bugs’ multifaceted and often external nature. We also found that non-linearity alone might not be the complete solution for Tensor bugs, as code structure analysis from information retrieval (BLUiR) helped to locate such structural bugs.
  • Complexities within the deep learning architectures: Deep learning systems present a unique set of challenges for bug localization due to their convoluted architecture. The complex nature of frameworks, characterized by multiple layers and components, often coupled with dependencies, hinders the performance of existing bug localization techniques. In contrast, the modular nature of libraries helps them in localization slightly by providing a more defined structure [57]. On the other hand, tools (e.g., TensorBoard) frequently integrate with larger frameworks (e.g., TensorFlow), which leads to dependencies and interactions, thus low performance in bug localization.
4.5 Implications
The above challenges highlight the necessity for novel bug localization techniques specifically designed for deep learning systems. Our research findings indicate that a one-size-fits-all approach may not be effective in practice, as shown by the existing methods (e.g., BLUiR, DNNLOC). Future research should focus on developing more comprehensive methods capable of addressing the wide array and complexity of bugs in deep learning systems, leveraging insights from the strengths and weaknesses of existing techniques identified in our research.

### v2

@@ 4.2 Answering RQ2: How do different types of bugs in deep learning systems impact bug localization?
For this example, DNNLOC retrieved the ground truth file at Top@4, while BLUiR ranked it at Top@7. bjXnet performs slightly better than DNNLOC in localizing training bugs and retrieving the ground truth file at Top@3. bjXnet’s inability to rank the correct file at Top@1 likely stems from the distributed nature of training bugs. These bugs often span across multiple interconnected components like callbacks, loss functions, and training loops, making it harder to pinpoint a single root cause. While bjXnet’s attention mechanism can highlight important sections like callbacks and loss scaling, training bugs tend to involve interactions between different parts of the code. This complexity means that the root cause may not be confined to a single location, requiring a broader contextual understanding of the training process, which limits bjXnet’s ability to achieve Top@1 localization
On the other hand, BLIA also performs poorly, ranking the correct file at Top@21. Meta information like stack traces and version history from BLIA might not be particularly helpful for training bugs, as they often involve complex, abstract issues (e.g., improper configurations of loss functions or callbacks), which do not leave clear traces in version control or stack traces. Moreover, BugLocator performs the worst (e.g., Top@23 position for the example bug in Table 24) in localizing the training bugs (Fig. 16) (as shown in Appendix A), which suggests that similarity analysis between bug reports and source code might not be sufficient for identifying these bugs. Training bugs, which occur during the model’s training phase and might involve aspects such as an improperly defined loss function, are more conceptual. These bugs might not be easily located using these existing methods and might require techniques that can offer a deeper insight into the model’s architecture and training process.
Tensor bugs: Tensors are central to deep learning and often involve intricate dimensional and mathematical issues. From Fig. 9, we notice that bjXnet, DNNLOC, and BLUiR are more effective in localizing tensor bugs than other techniques (e.g., BugLocator, BLIA). In the example bug from Table 25 [42], the bug report described the issue with the ‘nd.slice’ function in MXNet, which should return an empty tensor when the ‘begin’ and ‘end’ parameters are equal. Instead, the function returns a tensor with the same shape as the data.
@@
For this specific tensor bug, bjXnet, DNNLOC, and BLUiR all successfully retrieved the correct buggy file at Top@1 due to their targeted handling of tensor-related operations. bjXnet’s attention mechanism focuses on key operations, such as the slice function, which mishandles tensor shapes when begin equals end. This targeted focus might allow bjXnet to effectively identify the specific code causing the bug. DNNLOC captures the complex interactions between the bug report and tensor manipulations, mapping the error in how the nd.slice function handles dimensions directly to the buggy code. Interestingly, by parsing the AST of the code, BLUiR identified the relevant code snippet in the test_slice() function, which shares significant keyword overlap with the bug report. BLUiR determines the relevance of low-level code elements (e.g., class names, method names) against a bug report, which helps to reduce noise in the code segments where tensors could be handled or manipulated.
@@
API bugs: From Fig. 9, we observe that DNNLOC performs well in localizing API bugs, whereas BLUiR is the least effective. This could be due to BLUiR’s heavy reliance on structural information from the source code, which might not capture the specifics of API bugs (e.g., incorrect API calls). Another factor could be the rapid evolution of deep learning APIs, which affects versioning and compatibility [86]. Due to frequent structural changes in the code, the technique might be less effective in assessing the relevance between code and bug reports.
@@
Interestingly, despite its advanced attention mechanisms, bjXnet ranked the correct buggy file at the 6th position, lower than DNNLOC and BugLocator. We found that bjXnet’s attention mechanism might over-prioritize certain structural sections of the code, causing it to miss key interactions between components, which are critical for localizing API bugs. For example, in the unravel-index bug (Table 26), bjXnet correctly identified the key unravel-index function but failed to capture how the interactions between the API and different framework versions caused the bug. The issue arose from the shape parameter’s change in behavior across MXNet versions, but bjXnet’s focus on specific code segments may have led it to miss this broader context.
GPU bugs: Our investigation reveals that GPU bugs are the most difficult to localize for all five existing approaches. One possible explanation could be the complex nature of GPU bugs, as they can be triggered by a variety of factors, such as the compatibility between hardware (e.g., GPU device) and software (e.g., PyTorch). It might not even be located in the source code (a.k.a extrinsic GPU bug) [31]. From Table 29, we notice that only 17.65% of the GPU bugs can be found in the source code (a.k.a intrinsic GPU bug). Some examples of intrinsic GPU bugs that we found – the wrong reference to a GPU device, failed parallelism, incorrect state sharing between subprocesses, and faulty data transfer to a GPU device.
@@
Table 27 [44] shows a GPU bug triggered by the codebase (a.k.a intrinsic). The bug is connected to the use of multiple GPUs during training. According to the report, the machine contains multiple GPU devices, but only one GPU is used during computation. All five bug localization techniques performed poorly in locating the actual buggy code for this GPU bug. BLUiR retrieved the buggy file at the 50th position, followed by BugLocator at the 65th, BLIA at the 47th, DNNLOC at the 43rd, and bjXnet at the 42nd position in their respective rankings.
@@
We observed a minimal keyword overlap and structural similarity between the bug report (Table 27) and the actual buggy code (Fig. 20) (as shown in Appendix A), presenting a significant challenge in locating such bugs using BugLocator. Additionally, BLUiR’s analysis of smaller code segment similarity also proves ineffective for GPU bugs, as it solely concentrates on source code and overlooks the hardware-software interactions and runtime specifics crucial for comprehending such bugs. Similarly, BLIA’s integration of stack trace information, commit history, or version control history fails to encapsulate the unique aspects of GPU bugs as well. Moreover, despite DNNLOC’s and bjXnet’s capability to capture non-linear complex relationships through deep neural networks, it did not help to locate GPU bugs. One reason might be that these bugs often involve complex hardware-software dynamics that are not typically addressed by standard source code analysis or within the codebase’s non-linear mappings.
Table 28: Experimental result of existing bug localization techniques (BugLocator, BLUiR, BLIA, DNNLOC, bjXnet) for non-deep learning-related (NDL) bugs Bugs in deep learning systems
Method MAP MRR BugLocator 0.362 0.417 BLUiR 0.292 0.334 BLIA 0.437 0.381 DNNLOC 0.588 0.439 bjXnet 0.495 0.450
NDL bugs: We found that 35.20% of bugs in deep learning applications are not directly related to deep learning, but they impact system behavior (e.g., failed CI build due to GPU compatibility issues). These bugs are known as Non-Deep Learning-related (NDL) bugs in deep-learning systems. From Table 28, we find that the performance of existing techniques in localizing these bugs is also poor. DNNLOC outperforms other techniques, whereas BLUiR performs the lowest in locating NDL bugs. We observed that NDL bugs are less complex than their deep-learning counterparts. However, they are more prone to be extrinsic than the traditional bugs (48.15% from Table 29); we provided the details about extrinsic bugs in RQ3. Since existing baseline techniques focus on code-level artifacts only, they might fall short in detecting these bugs from deep learning systems.
@@
4.3 Answering RQ3: What are the implications of extrinsic bugs in deep learning systems for bug localization?
Most of the traditional bug localization techniques rely on the similarity between bug reports and source code [11, 6, 12, 17, 14, 13, 23, 53]. However, if a bug is of an extrinsic nature (e.g., originates from the operating system), simply relying on source code may not be effective for its localization.
@@
To investigate the impact of extrinsic bugs in deep learning systems, we performed another manual analysis following a set of steps as follows.
  • We conducted our manual analysis on a sample of 385 bug reports, which have a 95% confidence level and a 5% margin of error. This allows our findings to be generalizable to the original population of deep learning system bugs.
  • We manually labeled 385 bug reports from deep learning software (DLSW) and 385 from non-deep learning software (NDLSW) as either extrinsic or intrinsic. We selected equal amounts of bug reports using randomization principles to ensure a fair comparison during the evaluation of bug localization techniques. We used stratified random sampling [84] ensuring that our dataset was representative of both deep learning and traditional systems, reducing sampling bias that could affect the prevalence ratio of extrinsic and intrinsic bugs across categories.
  • We strictly adhered to established heuristics [37] to annotate extrinsic and intrinsic bugs. For example, a bug was classified as extrinsic only if there was clear evidence of external factors, such as system call failures or library version mismatches, documented in the bug report or related discussions.
  • We achieved a Cohen’s kappa [83] score of 0.87, indicating substantial agreement and reinforcing the reliability of our manual labeling process. We resolved disagreements through a systematic consensus-building process involving joint re-examination of disputed cases, guided by taxonomy definitions and cross-referenced with additional documentation to ensure accuracy and consistency.
  • We documented the manual analysis using an Excel sheet, which can be found in our replication package [66]. Each author spent a total of $\approx$20 hours on the analysis.
Prevalence ratio of extrinsic & intrinsic bugs: We found 40.00% extrinsic bugs within a total of 385 bug reports from deep learning systems (Denchmark dataset). The notion of extrinsic bugs is relatively new, especially in the case of bugs from deep learning systems. For a better comparison, we also manually inspected 385 bugs from non-deep learning systems (BuGL dataset) and determined the prevalence ratio of extrinsic and intrinsic bugs. We found only 10.65% extrinsic bugs in non-deep learning systems. Thus, deep learning systems contain almost four times more extrinsic bugs (Fig. 10) than non-deep learning systems.
@@
Prevalence ratio of extrinsic & intrinsic bugs from deep learning systems: We randomly select 100 samples for each type of bug from deep learning systems (same subsets from RQ2) and determined the prevalence ratio of extrinsic and intrinsic bugs for each type. Table 29 shows the results of our manual analysis for different bug categories in deep learning systems in terms of extrinsic and intrinsic bugs. We see that the prevalence ratios of deep learning-related extrinsic bugs range from 21.90% to 82.35%, whereas for non-deep learning-related bugs, the prevalence ratio is 48.15%. This suggests that the deep learning components of a software system might be more likely to trigger extrinsic bugs than non-deep learning components.
Localization of extrinsic & intrinsic bugs from both systems: To further analyze the impact of extrinsic bugs on bug localization, we experimented with our baseline techniques from RQ1 on extrinsic and intrinsic bugs separately. We chose 100 random bugs to evaluate the performance of all five techniques in bug localization from each category of both benchmark datasets. We repeated the evaluation three times using three different random subsets and then calculated the average result for a fair comparison.
Figure 11: Performance of existing approaches (BugLocator, BLUiR, BLIA, DNNLOC, bjXnet) for localizing extrinsic and intrinsic bugs in both systems (DLSW & NDLSW)
From Table 30, we notice that DNNLOC performs slightly better than other techniques in localizing extrinsic bugs for both systems. However, it shows a clear performance gap in localizing extrinsic and intrinsic bugs. The technique is less effective with extrinsic bugs, particularly in DLSW. It suggests a shortcoming in handling external complexities despite being able to capture non-linear relationships between bug reports and source code. We notice that BugLocator’s localization performance for extrinsic bugs in DLSW is lower than NDLSW. BugLocator might not be able to locate the extrinsic bugs in deep learning systems due to its naive approach, i.e., considering code as regular texts. We also found that BLUiR shows less performance gap between extrinsic and intrinsic bugs. It extracts different structured items (e.g., methods, classes) from the source code and bug reports. Thus, even if they reside outside the current codebase and are invoked from an external library, they could be matched with relevant keywords from a bug report.
Interestingly, BLIA performs slightly better for extrinsic bugs in DLSW than NDLSW, which could be possible due to its diverse use of meta components (e.g., stack traces, version control history). Stack traces from deep learning systems (DLSW) often contain more intricate information and dependencies (e.g., complex neural network data flows, dependencies on specialized libraries, GPU-related synchronization issues) related to the bugs, unlike stack traces from non-deep learning systems, which might not provide the same level of detailed information [88, 89].
@@
Method DLSW+EXT DLSW+INT NDLSW+EXT NDLSW+INT MAP BugLocator 0.287 0.359 0.375 0.405 BLUiR 0.277 0.369 0.301 0.359 BLIA 0.321 0.449 0.382 0.487 DNNLOC 0.334 0.486 0.390 0.505 bjXnet 0.310 0.510 0.367 0.470 MRR BugLocator 0.371 0.489 0.409 0.476 BLUiR 0.296 0.445 0.309 0.364 BLIA 0.385 0.518 0.404 0.505 DNNLOC 0.398 0.544 0.420 0.521 bjXnet 0.380 0.535 0.392 0.496 DLSW+EXT = Extrinsic Bug from Deep Learning Systems
@@
When considering bjXnet, its performance with extrinsic bugs is slightly lower than DNNLOC in DLSW but outperforms it in NDLSW. While bjXnet’s attention mechanism effectively isolates critical components in the code, extrinsic bugs often involve dependencies on external libraries and system-wide interactions that are harder to capture through localized attention. This limitation could explain bjXnet’s slightly reduced performance for extrinsic bugs, particularly in deep learning systems where external factors, like third-party libraries or hardware dependencies, play a significant role. Despite this, bjXnet’s ability to focus on key sections of the code still allows it to perform well with intrinsic bugs across both systems.
Overall, these results suggest that extrinsic bugs are hard to localize, whether related to deep learning or not. However, deep learning bugs with an extrinsic nature are the more difficult to localize. On the other hand, from Fig. 11, we note that the performance of all five approaches for intrinsic bugs in NDLSW is higher compared to the intrinsic bugs in DLSW, which supports the fact that the bugs related to deep learning algorithms (a.k.a DL bugs) from deep learning systems are more challenging to localize. Furthermore, we conducted appropriate statistical significance tests and found that the performance differences between extrinsic and intrinsic bugs in both systems are statistically significant, with a p-value of 0.014 for the difference between DLSW+EXT and DLSW+INT, and a p-value of 0.021 for NDLSW+EXT and NDLSW+INT, both indicating medium effect sizes, which reinforce the observed trends.
@@ 4.4 Key findings
Our empirical results reveal several key challenges that hinder bug localization in deep learning systems. All five baseline techniques (BugLocator [11], BLUiR [12], BLIA [23], DNNLOC [17], bjXnet [30]) performed significantly worse with deep learning systems than traditional software systems (e.g., $\sim$30lower MAP on average), confirming a substantial performance gap. Through our three research questions, we identified the following factors contributing to this performance gap:
1. Dominance of Extrinsic Bugs
Challenge: Deep learning systems contain a significantly higher number of extrinsic bugs. Extrinsic bugs are issues triggered by external factors such as hardware-software compatibility, operating systems, third-party libraries, or training data [37]. We found nearly four times more extrinsic bugs in deep learning systems than in traditional ones. Since these bugs stem from outside the source code (e.g., GPU driver incompatibilities), they often escape code-centric bug localization methods. All five baseline techniques from our study struggled to locate the extrinsic bugs. GPU bugs were among the hardest to localize, with all our techniques performing the worst in this category (e.g., MAP below 0.30 across all the baselines). The complex interplay of deep learning system layers (framework ¿ CUDA library ¿ GPU hardware) [91] makes their localization challenging. Using the Chi-Square test, we also found a strong statistical correlation (e.g., a p-value of $\sim$1.79e-14) between deep learning projects and extrinsic bugs.
Implication: Existing bug localization approaches mainly focus on source code analysis and thus miss critical extrinsic bugs. Appropriate tools and processes capturing external factors and environmental context will be needed to address extrinsic bugs in deep learning systems. Integrating keyword search or stack trace pattern analysis [92] into bug triage might help identify externally triggered failures. Similarly, categorizing extrinsic bugs into distinct classes – issues related to third-party libraries (e.g., version mismatches), hardware-related problems (e.g., GPU bugs), operating system incompatibilities, and dataset validity – can facilitate a more systematic localization process. For example, localizing hardware-specific extrinsic bugs may greatly benefit from targeted environmental checks, utilizing tools for GPU profiling [93] (e.g., NVIDIA Visual Profiler [^10] https://developer.nvidia.com/nvidia-visual-profiler and PyTorch Profiler [^11] https://pytorch.org/tutorials/recipes/recipes/profiler_recipe.html).
2. Limitations of Textual Representation
Challenge: Many existing techniques for bug localization (e.g., IR-based) heavily depend on textual similarity between bug reports and source code, which might not be sufficient for localizing complex, deep-learning bugs. For example, bugs related to gradient vanishing require a deeper conceptual understanding of neural network behaviour beyond what textual descriptions in bug reports typically convey (e.g., ‘the model is not learning’). Similarly, the bug reports could contain highly abstract issues (e.g., incorrect model architectures, incorrect loss function), which warrant the knowledge of the model architecture, hyperparameters, loss function, training procedure, and dataset. Our study revealed that existing techniques significantly struggled with such nuanced cases. Although the bug reports for deep learning systems might contain more code snippets compared to traditional software systems, these snippets did not substantially enhance bug localization, according to our findings.
Implication: Given the abstract concepts in bug reports and the complexities in many deep learning bugs, future localization methods or tools should carefully leverage domain-specific knowledge [6]. In particular, a comprehensive understanding of neural network concepts (e.g., layers, gradients, loss functions), common framework usage patterns, and data-processing pipelines should be injected into these tools. For example, the bug report stating ‘the model isn’t learning’ was traced back to the vanishing gradient problem caused by multiple sigmoid layers. A domain-aware technique might recognize this underlying connection and point toward these sigmoid layers. Another way to tackle the challenge is to incorporate rich contextual information into bug reports. Essential details (e.g., model configurations, hyperparameters, dataset characteristics, and runtime conditions) in the bug report might improve bug localization since they can provide critical information to map the errors with the root cause.
3. Dynamic Complexity during Training
Challenge: Bugs emerging during model training represent another major challenge. We identified training bugs as the most prevalent bug type ($\sim$27%) in deep learning systems. They involve dynamic behaviors such as model convergence and iterative parameter updates. Our baseline techniques often struggle to locate these bugs since they emerge from sequences of training steps and are not visible within the code. For example, a misconfigured batch size might only reveal itself after several epochs of training, rendering traditional, code-centric bug localization ineffective.
Implication: To effectively localize bugs manifesting during the model training phase, future techniques should incorporate dynamic analysis along with static analysis [7]. By capturing runtime information and associating it with the relevant source code, training-phase bugs, especially silent bugs, might be localized and diagnosed.
4. Lack of a Tailored Approach
Challenge: Our study indicates that no single technique performs consistently across all bug categories in deep learning systems. Techniques based on information retrieval (e.g., BLUiR) often perform well when textual cues in a bug report map directly to the source code (e.g., API bugs). On the other hand, deep learning–based approaches (e.g., DNNLOC, bjXnet) can more effectively address conceptual bugs (e.g., model bugs). However, neither approach fully addresses a major subset of bugs from deep learning systems, ranging from hardware compatibility problems to complex model-level inconsistencies, exposing the limitations of a ‘one-size-fits-all’ paradigm.
Implication: Unlike traditional software systems, different bug types in deep learning systems might benefit from different localization strategies. Future work should explore more holistic and adaptive techniques. Rather than relying on a single method, researchers could identify specific bug contexts (e.g., model, GPU, or API bug) and then apply appropriately tailored localization methods.